please help me answer this question:
a flare is launched from a life raft with an initial velocity of 80 meters per second. how many seconds will it take for the flare to return the sea?

Answer by stanbon(75887)

a flare is launched from a life raft with an initial velocity of 80 meters per second. how many seconds will it take for the flare to return the sea?
--------------
h(t) = -9.8t^2 + 80t
----
When the flare is at sea level, it height is zero.
---
Solve -9.8t^2+80t = 0
t(-9.8t+80) = 0
t= 0 when the flare leaves the raft.
-9.8t+80 = 0 when the flare hits the water.
9.8t = 80
t = 8.16 seconds
==========================
Cheers,
Stan H.